OpenStudy (anonymous):

how does cos((x+1)pi)/cos(x*pi)=-1?

jimthompson5910 (jim_thompson5910):

cos(n*pi) = 1 when n is an even integer cos(n*pi) = -1 when n is an odd integer

jimthompson5910 (jim_thompson5910):

so assuming x is an integer, then if x is odd, then x+1 is even (or vice versa)

jimthompson5910 (jim_thompson5910):

if x is odd, then x+1 is even making cos((x+1)pi) = 1 and cos(x*pi) = -1 ---------------------------- if x is even, then x+1 is odd making cos((x+1)pi) = -1 and cos(x*pi) = 1

jimthompson5910 (jim_thompson5910):

either way, they divide to -1

OpenStudy (anonymous):

cos(x+pi)=-cos(x)